# Estimating Bethe roots with VQE

### v1

@@
David Raveh111dxr921@miami.edu and  Rafael I. Nepomechie222nepomechie@miami.edu
@@ 2.1 Closed chain
The order of Bethe roots $k_{i}$ in $\vec{k}$ does not matter, since the Bethe equations are invariant under the interchange of any pair $(k_{j}\,,k_{l})$. Note also that the Bethe equations remain invariant under a $2\pi$ shift of any Bethe root, $k_{j}\mapsto k_{j}+2\pi$; hence, we can restrict $-\pi<\Re e(k_{j})\leq\pi$.
@@ 3 Ground-state Bethe roots
In this section, we apply the VQE algorithm, using Bethe states $|B^{L}_{M}(\vec{k})\rangle$ (2.5) as trial states and treating the Bethe roots $\vec{k}=(k_{1},\ldots,k_{M})$ as variational parameters, in order to estimate the Bethe roots corresponding to ground states. In this hybrid quantum/classical approach, where we aim to minimize the expectation value $\langle B^{L}_{M}(\vec{k})|\mathcal{H}|B^{L}_{M}(\vec{k})\rangle$, we evaluate the expectation values using a quantum simulator (ideally, a real device), using the algorithm in [15] to prepare the trial state $|B^{L}_{M}(\vec{k})\rangle$; and the variational parameters $\vec{k}$ are updated using a classical optimizer. We implement this approach using Qiskit v1.0 [18], with both the Statevector and Aer simulators, adopting cobyla as the classical optimizer.
We first consider in Sec. 3.1 the closed XXZ chain (2.1) with $L$ even in the antiferromagnetic regime $\Delta>0$, where the non-degenerate ground state is known to be described by $M=L/2$ real Bethe roots. We then consider in Sec. 3.2 the open XXZ chain (2.10), choosing values for the boundary fields such that the ground state is no longer described by only real Bethe roots.
@@ 3.1 Closed chain
We observe that the Statevector results are in excellent agreement with the true results. The Aer results are in fair agreement with the true results; increasing the number of shots improves the accuracy.
@@ 3.2 Open chain
We observe that the Statevector results are again in excellent agreement with the true results; and the Aer results are again in fair agreement with the true results.
@@ 4 Excited-state Bethe roots
$\langle\left(\mathcal{H}-\langle\mathcal{H}\rangle\right)^{2}\rangle=\langle B^{L}_{M}(\vec{k})|\mathcal{H}^{2}|B^{L}_{M}(\vec{k})\rangle-\langle B^{L}_{M}(\vec{k})|\mathcal{H}|B^{L}_{M}(\vec{k})\rangle^{2}\,,$ (4.1)
which is 0 for any exact eigenstate, see e.g. [17]. Since a 0 variance can also arise if $|B^{L}_{M}(\vec{k})\rangle$ is null, we always check that $\langle B^{L}_{M}(\vec{k})|\mathcal{H}|B^{L}_{M}(\vec{k})\rangle$ is not 0.555This check fails for the exceptional cases that the energy is zero, such as the case $(L,M)=(2,1)$ in Table 3.
It should be noted that this method makes it equally likely to obtain the Bethe roots of any particular eigenstate; that is, there is no bias in favor of low-energy eigenstates. Thus, obtaining the Bethe roots of a particular eigenstate requires randomly selecting initial parameter values $\vec{k}$ until the output roots correspond to the desired eigenstate.
@@ 4.1 Closed chain
Our results are displayed in Table 3. The columns are computed similarly as in Table 1. As for the ground-state Bethe roots, the Statevector results for excited-state Bethe roots are in excellent agreement with the true results, which confirms the viability of this variance-minimization approach. The Aer results are again in fair agreement with the true results.
@@ 4.2 Open chain
Our results are displayed in Table 4. The columns are computed similarly as in Table 1. Once again, we find excellent agreement between the Statevector and true results, and fair agreement between the Aer and true results.
@@ 5 Discussion
A limitation of this approach, which is also present in the classical approach of solving Bethe equations using Newton’s method, is the need for the initial values of the Bethe roots to be “sufficiently close” to the true values. As the number of real variational parameters (which is $\mathcal{O}(M)$) increases, this problem is exacerbated. A further limitation is the presence of shot noise, which however can be ameliorated by increasing the number of shots. On the other hand, an advantage of our VQE approach for ground-state Bethe roots over directly solving Bethe equations is that, in the former approach, the resulting Bethe roots are guaranteed to be “physical”; that is, they correspond to actual eigenstates of the Hamiltonian. (When directly solving Bethe equations, one often finds “unphysical” solutions, such as those containing repeated roots, which do not correspond to actual eigenstates, see e.g. [6].)

### v2

@@
David Raveh111dxr921@miami.edu,3 and  Rafael I. Nepomechie222nepomechie@miami.edu, [^3] Corresponding author
@@ 2.1 Closed chain
The order of Bethe roots $k_{i}$ in $\vec{k}$ does not matter, since the Bethe equations are invariant under the interchange of any pair $(k_{j}\,,k_{l})$. Note also that the Bethe equations remain invariant under a $2\pi$ shift of any Bethe root, $k_{j}\mapsto k_{j}+2\pi$; hence, we can restrict $-\pi<\Re e(k_{j})\leq\pi$. For $M=1$, the empty product on the right-hand-side of (2.3) is understood to give 1; the same applies to the empty products for $M=1$ in Eqs. (2.7), (2.11), (2.15), etc. below.
@@ 3 Ground-state Bethe roots
In this section, we apply the hybrid quantum/classical VQE algorithm [9], using Bethe states $|B^{L}_{M}(\vec{k})\rangle$ (2.5) as trial states and treating the Bethe roots $\vec{k}=(k_{1},\ldots,k_{M})$ as variational parameters, in order to estimate the Bethe roots corresponding to ground states. The key point is that, according to the variational theorem, for arbitrary values of $\vec{k}$ we have the inequality
$E_{0}\leq\langle B^{L}_{M}(\vec{k})|\mathcal{H}|B^{L}_{M}(\vec{k})\rangle\,,$ (3.1)
where the equality holds if and only if $\vec{k}$ are the Bethe roots corresponding to the ground state. The standard VQE algorithm begins by randomly selecting initial values $\vec{k}=\vec{k}^{(0)}$, and seeks to vary $\vec{k}$ so that the expectation (3.1) is minimized via the following steps:
  1. 1.
    Given $\vec{k}=\vec{k}^{(i)}$, prepare the state $|B^{L}_{M}(\vec{k}^{(i)})\rangle$ on a quantum computer.
  2. 2.
    Compute the expectation $\langle B^{L}_{M}(\vec{k}^{(i)})|\mathcal{H}|B^{L}_{M}(\vec{k}^{(i)})\rangle$ on a quantum computer.
  3. 3.
    Repeat steps 1-2 for some points $\vec{k}=\vec{k}^{(i)}+\vec{\epsilon}$ in an $\vec{\epsilon}$-neighborhood of $\vec{k}^{(i)}$. Use these values to compute with a classical optimizer the numerical gradient of $\langle B^{L}_{M}(\vec{k}^{(i)})|\mathcal{H}|B^{L}_{M}(\vec{k}^{(i)})\rangle$ about the point $\vec{k}^{(i)}$, and select $\vec{k}=\vec{k}^{(i+1)}$ in the direction of steepest descent.
  4. 4.
    Repeat steps 1-3 with $i\to i+1$ until the expectation converges to the ground-state energy.
We implement this approach using Qiskit v1.0 [18], which provides the necessary functionality for a given choice of trial state, Hamiltonian, and optimizer; our code is available as Supplementary Material. We simulate the results rather than using a real quantum device. In step 1, we prepare the state $|B^{L}_{M}(\vec{k}^{(i)})\rangle$ using the circuit in[15]. In step 2, the expectation $\langle B^{L}_{M}(\vec{k}^{(i)})|\mathcal{H}|B^{L}_{M}(\vec{k}^{(i)})\rangle$ is computed in one of two ways: using Qiskit’s Statevector simulator, which actually carries out the matrix multiplication (and hence is exact), or using Qiskit’s noiseless Aer simulator. The Aer simulator computes the expectation approximately, operating via shots, or repeated trials. [^4] There are several ways to measure the expectation of a Hamiltonian expressed as a sum $\mathcal{H}=\sum_{i}a_{i}\,U_{i}$ of Pauli strings $U_{i}$, e.g. $U_{i}=\sigma_{j}^{x}\sigma_{k}^{z}$ on a quantum computer. In one method, $p_{i}=\langle B^{L}_{M}(\vec{k})|U_{i}|B^{L}_{M}(\vec{k})\rangle$ can be computed via the Hadamard test, which requires repeated trials. It follows that $\langle B^{L}_{M}(\vec{k})|\mathcal{H}|B^{L}_{M}(\vec{k})\rangle=\sum_{i}a_{i}\,p_{i}$. For this algorithm to run on a real device, one would need to implement a shot-based computation, and so the Aer simulator results reflect the result of operating this algorithm on a perfect (noiseless) quantum computer. In step 3, we adopt Qiskit’s Cobyla as the classical optimizer. In step 4, we iterate the algorithm no more than 1000 times, although typically convergence occurs far sooner; and as the expectation converges to the ground-state energy, our variational parameters converge to the ground-state Bethe roots.
We first consider in Sec. 3.1 the closed XXZ chain (2.1) with $L$ even in the antiferromagnetic regime $\Delta>0$, where the non-degenerate ground state is known to be described by $M=L/2$ real Bethe roots. Next we consider in Sec. 3.2 the open XXZ chain (2.10), choosing values for the boundary fields such that the ground state is no longer described by only real Bethe roots. We then analyze the accuracy of the results in Sec. 3.3.
@@ 3.2 Open chain
3.3 Error
The Bethe root values displayed in Tables 1 and 2 are not of interest per se; rather, we focus on the error in our methodology, to verify its viability. One source of error is the imperfect classical optimizer, which of course behaves more poorly as $M$ increases; more sophisticated optimizers than Cobyla may exist that can perhaps improve upon this error. Additionally, the finite number of iterations in the VQE account for some error, which can be mitigated by increasing the number of iterations. An important source of error, unique to the Aer roots, is the error associated with the number of shots used to approximate an expectation value. This error $\epsilon$ scales as (see e.g. [20])
$\epsilon\sim\frac{1}{\sqrt{x}}\,,$ (3.2)
where $x$ is the number of shots. Thus, with sufficiently many shots this error can be mitigated up to error $\epsilon$ for fixed $L$; the number of expectation values computed in a single VQE iteration grows with $L$, see Footnote 4. On a real device, there would of course be additional errors due to various types of noise.
We observe in Tables 1 and 2 that the Statevector results, which represents a VQE with zero shot-based error, are in excellent agreement with the true results. The Aer results, which represent the result of running the VQE algorithm on a noiseless quantum computer, are in fair agreement with the true results; increasing the number of shots improves the accuracy according to (3.2), see Fig. 1.
Figure 1: The error of the Aer roots $\vec{k}^{\prime}$ with respect to the True roots $\vec{k}$, defined as $\epsilon=\|\vec{k}-\vec{k}^{\prime}\|/\|\vec{k}\|$, where $\|\vec{k}\|=\sqrt{\vec{k}\cdot\vec{k}^{*}}$, for the ground-state Bethe roots of the closed (blue) and open (red) chains when $L=4$, with curves fitted to $\epsilon=a/\sqrt{x}+b$ (3.2).
@@ 4 Excited-state Bethe roots
$0\leq\langle\left(\mathcal{H}-\langle\mathcal{H}\rangle\right)^{2}\rangle=\langle B^{L}_{M}(\vec{k})|\mathcal{H}^{2}|B^{L}_{M}(\vec{k})\rangle-\langle B^{L}_{M}(\vec{k})|\mathcal{H}|B^{L}_{M}(\vec{k})\rangle^{2}\,,$ (4.1)
where the equality holds for any exact eigenstate, see e.g. [17].
It should be noted that this method makes it equally likely to obtain the Bethe roots of any particular eigenstate; that is, there is no bias in favor of low-energy eigenstates. Thus, obtaining the Bethe roots corresponding to a particular eigenstate requires randomly selecting initial parameter values $\vec{k}^{(0)}$ until the output roots correspond to the desired eigenstate.
@@ 4.1 Closed chain
Our results are displayed in Table 3. The columns are computed similarly as in Table 1. As for the ground-state Bethe roots, the Statevector results for excited-state Bethe roots are in excellent agreement with the true results, which confirms the viability of this variance-minimization approach. The Aer results are again in fair agreement with the true results; the error behaves similarly as described in Section 3.3.
@@ 4.2 Open chain
Our results are displayed in Table 4. The columns are computed similarly as in Table 1. Once again, we find excellent agreement between the Statevector and true results, and fair agreement between the Aer and true results, similar to the description in Section 3.3.
@@ 5 Discussion
A limitation of this approach, which is also present in the classical approach of solving Bethe equations using Newton’s method, is the need for the initial values of the Bethe roots $\vec{k}^{(0)}$ to be “sufficiently close” to the true values. As the number of real variational parameters (which is $\mathcal{O}(M)$) increases, this problem is exacerbated. A further limitation is the presence of shot noise, which however can be ameliorated by increasing the number of shots, see Figure 1. On the other hand, an advantage of our VQE approach for ground-state Bethe roots over directly solving Bethe equations is that, in the former approach, the resulting Bethe roots are guaranteed to be “physical”; that is, they correspond to actual eigenstates of the Hamiltonian. (When directly solving Bethe equations, one often finds “unphysical” solutions, such as those containing repeated roots, which do not correspond to actual eigenstates, see e.g. [6].)